John is 13 years less than 3 times as old as his son. Six years ago, he was 14 years more than twice as old as his son was then. How old is John?
Answer by ptaylor(2198)

Let x=son's age now; x-6=son's age six years ago
Then 3x-13=John's age now;3x-13-6=3x-19=John's age 6 years ago
Now we are told that:
3x-19=2(x-6)+14 or
3x-19=2x-12+14
3x-19=2x+2 add 19 to each side and subtract 2x from each side
3x-2x-19+19=2x-2x+2+19
x=21 son's age now
3x-13=3*21-13=50 John's age now
CK
3*21-19=2*21+2
44=44
Hope this helps--ptaylor